# It’s Not a Modality Gap: Characterizing and Addressing the Contrastive Gap

### v1

@@ 5 Experiments
  • $L_{\text{CLIP+Uniform+Align}}$ ($L_{\text{CUA}}$ ): $L_{\text{CLIP}}$ + $L_{\text{Align}}$ + $L_{\text{Uniform}}$
  • $L_{\text{CLIP+Uniform+Align+XUniform}}$ ($L_{\text{CUAXU}}$ ): $L_{\text{CLIP}}$ + $L_{\text{Align}}$ + $L_{\text{Uniform}}$ + $L_{\text{XUniform}}$
@@ 5.1 Effects of Reducing the Contrastive Gap on MS COCO
Finally, we evaluated the quality of the representational space after reducing the contrastive gap. We used the image-text retrieval task on the MS COCO validation dataset, a standard representative task in the vision-language domain. We present the results in Table 4. For the text-retrieval task ($I\rightarrow T$), all models have similar performance, with fine-tuned CLIP outperforming by a small margin. This implies that the text-retrieval task is not well correlated with uniformity, alignment, and size of contrastive gap. However, for the image-retrieval task ($T\rightarrow I$), we see that the models trained with uniformity and alignment terms better. Thus, image-retrieval likely benefits from increased uniformity and alignment, which leads to a smaller contrastive gap.
@@
Table 4: Image to text ($I\rightarrow T$) and text to image ($T\rightarrow I$) retrieval accuracies for MS COCO. Training with uniformity/alignment maintains image-text retrieval accuracies compared to default CLIP. We report 1-standard error confidence intervals in Appendix 10
@@ 5.2 Zero-Shot Transfer
We evaluate our fine-tuned CLIP models on five standard image-classification datasets: ImageNet1k (Russakovsky et al. (2015)), CIFAR 10, CIFAR 100 (Krizhevsky et al. (2009)), Caltech101 (Li et al. (2022)), and Describable Textures Dataset (DTD) (Cimpoi et al. (2014)). We adopt the evaluation strategy of Goel et al. (2022) and recommended by Radford et al. (2021): generate prompts using the class names to form sentences like "a photo of a class name", "A sketch of a "class name", etc. We then pass these sentences through the text encoder to get prompt embeddings. We average all prompt embeddings to a class embedding. We then pass the image through the image encoder and classify the image by finding the closest class via cosine similarity between the embeddings.
Figure 3 shows the average zero-shot accuracies across the five datasets for each of the losses and dimensionalities. CLIP losses with alignment/uniformity consistently outperform the default CLIP loss, and XUniform adds additional benefit. In Table 5, we present centroid distances between class and image embeddings, as well as uniformity loss values for image embeddings. We observe that $L_{\text{CUA}}$ and $L_{\text{CUAXU}}$ are able to learn representations with higher uniformity and smaller contrastive gap. $L_{\text{CUAXU}}$ performs the best across all the datasets and dimensionalities in the zero-shot classification task, suggesting that lower $L_{\text{Uniform}}^{I}$ and lower values for centroid distance (lower contrastive gap) may lead to better representations learned.
@@ B.3 Model Architecture and Hyperparameters
Text Encoder Transformer (same as in 4443)

### v3

@@ 5 Experiments
  • $L_{\text{CLIP+Uniform+Align}}$ ($L_{\text{CUA}}$ ): $L_{\text{CLIP}}$ + $L_{\text{Uniform}}$ + $L_{\text{Align}}$
  • $L_{\text{CLIP+Uniform+Align+XUniform}}$ ($L_{\text{CUAXU}}$ ): $L_{\text{CLIP}}$ + $L_{\text{Uniform}}$ + $L_{\text{Align}}$ + $L_{\text{XUniform}}$
@@ 5.1 Effects of Reducing the Contrastive Gap on MS COCO
Finally, we evaluated the quality of the representational space after reducing the contrastive gap. We used the image-text retrieval task on the MS COCO validation dataset, a standard representative task in the vision-language domain. We present the results in Table 4. For both the image-retrieval ($T\rightarrow I$) and the text-retrieval ($I\rightarrow T$) tasks, we see that models fine-tuned with all the loss functions perform comparably. These results suggest that it is possible to shrink the gap between the embeddings, while maintaining similar image-text retrieval performance after fine-tuning.
@@
Table 4: Image to text ($I\rightarrow T$) and text to image ($T\rightarrow I$) retrieval accuracies for MS COCO. Training with uniformity/alignment maintains image-text retrieval accuracies compared to default CLIP. We report 1-standard error confidence intervals in Table 10 (Appendix A.2).
@@ 5.2 Zero-Shot Transfer
We evaluate our fine-tuned CLIP models on five standard image-classification datasets: ImageNet1k (Russakovsky et al. (2015)), CIFAR 10, CIFAR 100 (Krizhevsky et al. (2009)), Caltech101 (Li et al. (2022)), and Describable Textures Dataset (DTD) (Cimpoi et al. (2014)). We adopt the evaluation strategy of Goel et al. (2022) and recommended by Radford et al. (2021): We generate prompts using the class names to form sentences like "a photo of a class name", "A sketch of a "class name", etc. We then pass these sentences through the text encoder to get prompt embeddings. We average all prompt embeddings to a class embedding. We then pass the image through the image encoder and classify the image by finding the closest class via cosine similarity between the embeddings.
Figure 3 shows the average zero-shot accuracies across the five datasets for each of the losses and dimensionalities. CLIP losses with alignment/uniformity consistently outperform the default CLIP loss, and XUniform adds additional benefit. In Table 5, we present centroid distances between class and image embeddings, as well as uniformity loss values for image embeddings. We observe that $L_{\text{CUA}}$ and $L_{\text{CUAXU}}$ are able to learn representations with higher uniformity and smaller contrastive gap. $L_{\text{CUAXU}}$ performs the best on average across all the datasets and dimensionalities in the zero-shot classification task, suggesting that lower $L_{\text{Uniform}}^{I}$ and lower values for centroid distance (lower contrastive gap) may lead to better representations learned.
@@ A.1 Gap Metrics on MS COCO after fine-tuning
Table 7: Gap metrics on MS COCO validation dataset for CLIP in $\mathbb{R}^{32}$.
@@ B.3 Model Architecture and Hyperparameters
Text Encoder Transformer (same as in footnotemark: )